A sample of O2 gas with a volume of 3.03 L was collected at 272 °C. What is the volume of the gas if the temperature is changed to 106 °C at constant pressure
1
Expert's answer
2017-02-24T14:47:05-0500
Here we apply the Charles’s law which states that temperature of gas and its volume are directly related in case of constant pressure of the sample:
V / T = k,
where
V – volume of gas, L;
T – temperature of gas, K;
k = constant.
If temperature changes from T0 to T1 and respective volume from V0 to V1 than we have:
V0 / T0 = V1 / T1,
And therefore
V1 = V0 * T1 / T0.
To convert temperature from °C to K we have to add 273.15.
Do the calculation:
V1 = 3.03 L * (106 + 273.15) K / (272 + 273.15) K = 2.11 L
Answer:
The volume of the gas is 2.11 L.